A small particle has charge -5.40 μC and mass 3.00×10−4 kg. It moves from point A, where the electric potential is VA = 240 V, to point B, where the electric potential VB = 810 V is greater than the potential at point A. The electric force is the only force acting on the particle. The particle has a speed of 6.00 m/s at point A.
Answer #1

from Conservation of energy change in kol. = change in Ilmu ² - 1 mu² = (VB-VA) q P.E. m. - v² = u²+ 29 (UG-VA) v2 = 36 + 2X5hope this helps.

Please rate positively.

Please comment for any doubt or further clarification...